OpenStudy (anonymous):

The sum of two numbers is 12. If one number is subtracted from the other, the result is 0. Find the numbers.

OpenStudy (jdoe0001):

a-b = 0 if you were to solve for "a", what would that look like?

OpenStudy (anonymous):

-b=0-a

OpenStudy (jdoe0001):

and if you multiply both sides by -1, what would that look like?

OpenStudy (anonymous):

-1(-b)=(-1)0-a

OpenStudy (jdoe0001):

ok, and if we expand that, what's -1 * -b?

OpenStudy (anonymous):

1b

OpenStudy (jdoe0001):

so -1 * -a will be 1a 1a = a 1b = b thus b = a

OpenStudy (jdoe0001):

so we know about this 2 numbers is that a + b = 12 a - b = 0 we also know now that, b = a so if I were to rewrite the 1st equation with b = a in mind, it'd look like a + b = 12 OR a + (a) = 12 a + a = 2a, thus 2a = 12 what if we divide both sides by 2?

OpenStudy (anonymous):

a=6

OpenStudy (jdoe0001):

hmm, low and behold! a = 6 so if I were to use that in the 1st equation let's see a + b = 12 OR 6 + b = 12 # substracting 6 from each side b = 12 -6 =6 b = 6

OpenStudy (jdoe0001):

well, keep in mind that we knew that b = a, if a = 6, then b = 6 too anyway :)

OpenStudy (jdoe0001):

so let's check that a + b = 12 a - b = 0 6 + 6 = 12 6 - 6 = 0

OpenStudy (anonymous):

....you made that so much easier than i thought it was. ugh. thank you! :)

OpenStudy (jdoe0001):

yw